# Lipschitz-Guided Design of Interpolation Schedules in Generative Models

### v1

@@ Abstract.
We study the design of interpolation schedules in the stochastic interpolants framework for flow and diffusion-based generative models. We show that while all scalar interpolation schedules achieve identical statistical efficiency under Kullback-Leibler divergence in path space after optimal diffusion coefficient tuning, their numerical efficiency can differ substantially. This observation motivates focusing on numerical properties of the resulting drift fields rather than statistical criteria for schedule design. We propose averaged squared Lipschitzness minimization as a principled criterion for numerical optimization, providing an alternative to kinetic energy minimization used in optimal transport approaches. A transfer formula is derived that enables conversion between different schedules at inference time without retraining neural networks. For Gaussian distributions, our optimized schedules achieve exponential improvements in Lipschitz constants over standard linear schedules, while for Gaussian mixtures, they reduce mode collapse in few-step sampling. We also validate our approach on high-dimensional invariant distributions from stochastic Allen-Cahn equations and Navier-Stokes equations, demonstrating robust performance improvements across resolutions.
@@ 1.1. Context
Dynamics between probability measures, particularly flows and diffusion processes described by ordinary and stochastic differential equations (ODEs and SDEs), form the foundation of state-of-the-art generative modeling techniques [38, 19, 42]. These methods generate samples through an iterative refinement process that progressively eliminates noise or corruption at different scales [40, 23].
In this paper, we study the impact and design of interpolation schedules on the performance of flow and diffusion-based generative models. We work within the stochastic interpolant framework [1, 2], which provides a systematic approach for modeling noising processes through sample interpolation and enables principled construction of the corresponding generative processes. This framework connects to related concurrent work on flow matching [27] and rectified flows [28], and encompasses diffusion and score-based generative models [38, 42, 19, 41] as specific instances.
@@ 1.2. Basics of stochastic interpolants
For different values of $t$, the interpolant $I_{t}$ can be seen as modeling a corruption of the target at a specific scale. The theory of stochastic interpolants [2, 1] shows that one can generate samples from $\mu^{*}$ by solving the following ODE:
@@
where $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$ and $\dot{I}_{t}$ denotes the time derivative of $I_{t}$. The solution satisfies $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$, and in particular, $X_{1}\sim\mu^{*}$. This can also be seen as a consequence of the mimicking theorem [17], also referred to as Markovian projection.
@@ 1.3. This work
  • In Section 2, we prove that under the Kullback-Leibler divergence criterion in path space, different choices of scalar schedules are statistically equivalent when diffusion coefficients are optimized a posteriori. This equivalence paradoxically renders statistical considerations insufficient for schedule selection.
@@ 1.4. Related work
Since the introduction of flow and diffusion models, numerous studies have examined the design principles and parameter space of these models (see a review in [48]). These investigations encompass the choice of noise, noising processes, time reversal processes, training losses, and diffusion coefficients in the generative process. This paper focuses on designing interpolation schedules within the framework of unit-time stochastic interpolants [2, 1], which relates to the noising and denoising schedules in diffusion models.
@@
Schedule design is of interest from both statistical and numerical perspectives. From a statistical standpoint, it was demonstrated in [25] that different noise schedules in diffusion models yield the same variational lower bound. Our results suggest that this “statistical equivalence” generalizes to a broader context using the unit-time stochastic interpolants framework, with the Kullback-Leibler divergence in path space serving as the statistical estimation criterion (see discussion in Remark 2.6).
From a numerical perspective, existing works have derived insights primarily through empirical studies on machine learning datasets to tune the noise schedules for efficient sampling performance [35, 22, 30, 41, 23]; see also [36, 47, 32, 5] for learning improved schedules with additional training and an analysis [45] considering score errors. In this work, we propose a principled way for numerical design by optimizing the Lipschitzness of the drift field at inference time. Related mathematical work focused on the Lipschitz regularity of flows and flow maps includes [9, 44]. See also [3] for a mathematical investigation of the numerical impact of schedules on identifying modes in high dimensions.
It has been advocated to learn the optimal transport path [28], which is straight and therefore offers better numerical performance; see also generative models built using entropy-regularized optimal transport, namely Schrödinger bridges [11, 37]. Nevertheless, the optimal transport path may lead to irregular drift fields [44] that are not ideal for numerical integration (see also an example in Remark 3.4), an issue our proposed criterion of optimizing the Lipschitzness aims to address.
Moreover, there has been a line of work on improving numerical performance with high order, exponential, or parallel integrators, e.g., [13, 29, 50, 26, 6, 46, 10, 43], and multiscale and cascading approach [49, 12, 21, 33, 20, 16, 31], which can be combined with the design of schedules to accelerate sampling. We also note another line of work focused on consistency models and learning flow maps (e.g., [39, 24, 34, 14, 4]), to achieve few steps sampling and thus improve numerical efficiency.
@@ 2.1. Stochastic interpolants
Here we briefly recall the main results of the stochastic interpolant framework [2, 1]. For completeness, we also include a simple sketch of derivations in Appendix A.
@@ 2.3. Optimizing the KL in path space
Given the flexibility of choosing $\epsilon_{t}$, it is natural to ask which $\epsilon_{t}$ is optimal. Let us consider the criterion of the KL divergence between path measures $\mathbb{P}_{X}$ and $\mathbb{P}_{\hat{X}}$ of $X=(X_{t})_{0\leq t\leq 1}$ and $\hat{X}=(\hat{X}_{t})_{0\leq t\leq 1}$, respectively. Aaccording to Girsanov’s theorem, this KL divergence has the form
(2.5) $\mathrm{KL}[\mathbb{P}_{X}\|\mathbb{P}_{\hat{X}}]=\frac{1}{2\epsilon_{t}}\int_{0}^{1}\left(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t}\right)^{2}\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t\,.$
@@ 2.4. Equivalence over scalar schedules
In [25], it was pointed out that in diffusion models, different noise schedules lead to the same variational lower bound. In the continuous setting, this corresponds to the KL divergence in path space. Our results generalize their discussion to stochastic interpolants and incorporate the step of a posteriori tuning of diffusion coefficients.  $\lozenge$
Proposition 2.5 shows that the optimal KL accuracy in path space depends solely on the estimation of $\nabla\log q_{r}(x_{1}+rz)$: that is, from the perspective of KL divergence in path space, all linear scalar interpolants with independently coupled endpoints and one endpoint Gaussian are statistically indistinguishable. This indicates that other metrics need to be explored if we want to select models for improved statistical efficiency. On the other hand, using matrix-valued instead of scalar schedules may potentially lead to different statistical efficiency, a direction of interest in future work.
@@ 3. Numerical Design by Optimizing Averaged Squared Lipschitzness
The discussion in the previous section does not consider numerical efficiency: while different scalar schedules are statistically equivalent, they lead to ODEs or SDEs with dramatically different regularity properties of the drift term. In this section, we explore how to choose interpolation schedules that enhance numerical efficiency. We focus on ODEs rather than SDEs for simplicity, noting that ODEs typically achieve better empirical performance due to their greater ease of integration [23, 13].
@@ 3.1. From one schedule to another
The proposition implies that we can easily change the interpolation schedule from one to another if we know the true drift functions. This also applies to the estimators of the drift functions, so we can tune the schedule at inference time rather than during training. Similar statements have appeared in the literature [25, 23]. We will use this fact in numerical experiments in Section 4. The natural question now is which schedule to choose in practice.
@@ 3.2. Optimizing averaged squared Lipschitzness
The time-dilated schedule studied in [3] also resolve the mode missing issue:
@@ 3.5. High dimensional examples
This can be proved using the Cramér–Rao and Brascamp–Lieb inequalities; see [15]. Therefore, similar to the Gaussian case, we can choose $\lambda^{\star}=\lambda_{M}^{-1}$. Then, with the schedule
@@ 4. Numerical Demonstrations
In this section, we conduct numerical experiments to demonstrate the improved efficiency of the schedule designed in the previous section. We first study high-dimensional Gaussian distributions and Gaussian mixtures, followed by the high-dimensional invariant distributions of the stochastic Allen-Cahn equation and the Navier-Stokes equations, which exhibit slightly and highly non-Gaussian behaviors respectively.
We use the UNet architecture popularized by [19] to train the drift field for all experiments except Gaussian and mixtures, where we employ explicit formula. In all the examples, we integrate the ODE from $t_{\min}=10^{-3}$ to $t_{\max}=1-10^{-3}$ to avoid potential numerical issues at $t=0$ or $1$. Code is available at https://github.com/yifanc96/GenerativeDynamics-NumericalDesign.git.
For accuracy evaluation, we use the energy spectrum (or enstrophy spectrum in the case of Navier-Stokes) of the samples as the criterion. The spectrum for a sample $u$ (which is a function that is either 1D or 2D in this paper) is computed using the formula
$E(k)=\sum_{k\leq|m|_{2}\leq k+1}|\hat{u}(m)|^{2}\,$,
where $\hat{u}(m)$ are the Fourier coefficients. We average $E(k)$ over sufficiently many samples for each frequency $k$.
@@ 4.1. Gaussians
We consider the Gaussian random field $\mathsf{N}(0,\sigma^{2}(-\Delta+\tau^{2}I)^{-s})$, where $\Delta$ is the negative Laplacian with homogeneous Dirichlet boundary conditions on $D=[0,1]^{2}$. The true data distribution $x_{1}$ is sampled from this distribution with parameters $s=3$, $\tau=1$, and $\sigma^{2}=(4\pi^{2}+\tau^{2})^{s}$. The noise $z$ in the interpolant is sampled from white noise (which corresponds to $\sigma=1,s=0$ in the Gaussian random field). We discretize the 2D field on a grid with $N$ points in each dimension and construct flow-based generative models. The ODE is solved using the fourth-order Runge-Kutta (RK4) scheme.
Figure 2 shows random fields generated using a standard linear schedule $\beta_{t}=t$ compared to those using our designed schedule (3.11) optimized for avg-Lip2, both employing 20 RK4 steps with $N=128$. The designed schedule clearly produces superior samples. The right panel of Figure 2 displays both schedules: notably, the designed schedule exhibits rapid initial growth.
In Figure 3, we compare the energy spectra of the true distribution and generated samples. The designed schedule yields a more accurate spectrum, and this accuracy remains robust as resolution increases, unlike standard linear schedules where performance degrades with refinement.
@@ 4.2. Gaussian mixtures
We consider the $d$-dimensional Gaussian mixture distribution in (3.13) with $d=1000$, $p=0.3$, and $r=[1,1,\ldots,1]\in\mathbb{R}^{d}$. The noise $z$ is sampled from $\mathsf{N}(0,\mathrm{I})$. We compare the linear schedule $\beta_{t}=t$ and the approximate min-avg-Lip2 schedule (3.14); in both cases, $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$. We use the explicit formula for the drift given in Example 3.12 and run only $2$, 33, or $4$ steps of RK4 to integrate the ODE with $10^{4}$ independent noise samples. For the obtained samples, we use PCA to obtain a 1D projection and fit a 1D bimodal Gaussian mixture model to estimate the weights of the two modes.
@@
In Table 1, we compare results using the linear schedule and the approximate min-avg-Lip2 schedule. The latter clearly achieves better accuracy, while the former is prone to missing modes.

### v2

@@ Abstract.
We study the design of interpolation schedules in the stochastic interpolants framework for flow and diffusion-based generative models. We show that while all scalar interpolation schedules achieve identical statistical efficiency under Kullback-Leibler divergence in path space after optimal diffusion coefficient tuning, their numerical efficiency can differ substantially. This motivates focusing on numerical properties of the resulting drift fields rather than purely statistical criteria for schedule design. We propose averaged squared Lipschitzness minimization as a principled criterion for numerical optimization, providing an alternative to kinetic energy minimization used in optimal transport approaches. A transfer formula is derived that enables conversion between different schedules at inference time without retraining neural networks. For Gaussian distributions, the optimized schedules achieve exponential improvements in Lipschitz constants over standard linear schedules, while for Gaussian mixtures, they reduce mode collapse in few-step sampling. We also validate our approach on high-dimensional invariant distributions from stochastic Allen-Cahn equations and Navier-Stokes equations, demonstrating robust performance improvements across resolutions.
@@ 1.1. Context
Dynamics between probability measures, particularly flows and diffusion processes described by ordinary and stochastic differential equations (ODEs and SDEs), underpin many modern generative modeling techniques [sohl2015deep, ho2020denoising, song2020score]. These models generate samples by progressively transforming simple noise into structured data through a sequence of intermediate distributions, implemented as an iterative denoising or refinement process [song2019generative, karras2022elucidating]. The manner in which these transformations are traversed in time—encoded by interpolation or noise schedules—plays a crucial role in shaping the resulting generative dynamics and their performance.
In this paper, we study the mathematical design of interpolation schedules for flow and diffusion-based generative models within the stochastic interpolants framework [albergo2023stochastic, albergo2022building]. This framework provides a unified formulation for noising and denoising processes based on sample interpolation, and enables principled construction of the associated generative dynamics. It is closely related to concurrent developments such as flow matching [lipman2022flow] and rectified flows [liu2022flow], and encompasses diffusion and score-based generative models [sohl2015deep, song2020score, ho2020denoising, song2020improved] as special cases.
@@ 1.2. Basics of stochastic interpolants
For different values of $t$, the interpolant $I_{t}$ can be seen as modeling a corruption of the target at a specific scale. The theory of stochastic interpolants [albergo2022building, albergo2023stochastic] shows that one can generate samples from $\mu^{*}$ by solving the following ODE:
@@
where $b_{t}(x)=\mathbb{E}[\dot{I}_{t}|I_{t}=x]$ and $\dot{I}_{t}$ denotes the time derivative of $I_{t}$. The solution satisfies $\mathrm{Law}(X_{t})=\mathrm{Law}(I_{t})$, and in particular, $X_{1}\sim\mu^{*}$. This can also be seen as a consequence of the mimicking theorem [gyongy1986mimicking], also referred to as Markovian projection.
@@ 1.3. This work
  • In Section 2, we prove that under the Kullback-Leibler divergence criterion in path space, different choices of scalar schedules are statistically equivalent when diffusion coefficients are optimized a posteriori. This equivalence fundamentally renders statistical considerations insufficient for schedule selection.
@@ 1.4. Related work
Since the introduction of flow and diffusion-based generative models, a large body of work has explored their design principles and parameter choices; see the survey [yang2023diffusion]. These studies address, among other aspects, the choice of noise distributions, noising and denoising processes, time-reversal dynamics, training objectives, and diffusion coefficients. The present work focuses on the design of interpolation schedules within the unit-time stochastic interpolants framework [albergo2022building, albergo2023stochastic], which relates to concurrent developments in flow matching [lipman2022flow] and rectified flows [liu2022flow], and encompasses diffusion and score-based generative models [sohl2015deep, song2020score, ho2020denoising, song2020improved] as special cases. Within this framework, interpolation schedules play a role analogous to noise schedules in diffusion models.
@@
Schedule design has been studied from both statistical and numerical perspectives. From a statistical standpoint, [kingma2021variational] showed that different noise schedules in diffusion models yield the same variational lower bound. Our results indicate that this notion of statistical equivalence extends to a broader setting within the unit-time stochastic interpolants framework when the Kullback–Leibler divergence in path space is used as the estimation criterion; see Remark 2.6. This observation suggests that statistical considerations alone are insufficient to distinguish among scalar interpolation schedules.
From a numerical perspective, most existing work has relied on empirical studies—primarily on machine learning benchmarks—to tune noise or time schedules for improved sampling efficiency [san2021noise, jolicoeur2021gotta, nichol2021improved, song2020improved, karras2022elucidating]. Related approaches learn improved schedules or time parametrizations using additional training [shaul2024bespoke, xue2024accelerating, sabouralign, chentrajectory], while [wang2024evaluating] analyzes the sensitivity of schedule choice to score estimation errors. Here we propose a principled approach to numerical schedule design based on optimizing the Lipschitz regularity of the drift field at inference time, without requiring retraining.
Related mathematical work has investigated Lipschitz regularity, contractivity, and stability properties of flows and flow maps [daniels2025contractivity, tsimpos2025optimal]. In particular, [tsimpos2025optimal] considers a variational problem for minimizing the maximum Lipschitz constant under reparametrization of a fixed flow, a setting which differs from ours. See also [aranguri2025optimizing] for a mathematical study of how schedule design affects mode identification in high-dimensional distributions.
Another closely related line of work advocates learning optimal transport paths [liu2022flow, pooladian2023multisample], which are straight and therefore appealing from a numerical standpoint; related models based on entropy-regularized optimal transport include Schrödinger bridges [de2021diffusion, shi2023diffusion, pooladian2025plug]. However, optimal transport paths may give rise to highly irregular drift fields [tsimpos2025optimal], which are not well suited for numerical integration (see also Remark 3.4). Moreover, since the estimation of the initial drift $b_{0}$ inevitably contains errors, one-step or few-step generation strategies that rely heavily on straightness may struggle to accurately reproduce fine-scale features. Our proposed Lipschitz-based criterion is instead designed to directly target numerical integration efficiency and robustness to discretization.
Finally, numerical efficiency can also be improved through advances orthogonal to schedule design, including higher-order, exponential, and parallel integrators [dockhorn2022genie, lu2022dpm, zhang2022fast, li2024accelerating, chen2024accelerating, wu2024stochastic, de2025accelerated, tan2025stork], as well as multiscale and cascading approaches [yu2020wavelet, dhariwal2021diffusion, jing2022subspace, saharia2022image, ho2022cascaded, guth2022wavelet, phung2023wavelet]. In addition, consistency models and approaches based on learning flow maps [song2023consistency, kim2024consistency, salimans2022progressive, frans2024one, boffi2025flow] aim to reduce the number of sampling steps altogether. These approaches are complementary to schedule design and can be combined with the methods studied in this work.
We note that interpolation schedules can also substantially influence training stability and efficiency. The present paper focuses primarily on the statistical and numerical efficiency of schedule design.
@@ 2.1. Stochastic interpolants
Here we briefly recall the main results of the stochastic interpolant framework [albergo2022building, albergo2023stochastic]. For completeness, we also include a simple sketch of derivations in Appendix A.
@@ 2.3. Optimizing the KL in path space
Given the flexibility of choosing $\epsilon_{t}$, it is natural to ask which $\epsilon_{t}$ is optimal. Let us consider the criterion of the KL divergence between path measures $\mathbb{P}_{X}$ and $\mathbb{P}_{\hat{X}}$ of $X=(X_{t})_{0\leq t\leq 1}$ and $\hat{X}=(\hat{X}_{t})_{0\leq t\leq 1}$, respectively. According to Girsanov’s theorem, this KL divergence has the form
(2.5) $\mathrm{KL}[\mathbb{P}_{X}\|\mathbb{P}_{\hat{X}}]=\frac{1}{2\epsilon_{t}}\int_{\mathbb{R}^{d}}\int_{0}^{1}\left(\alpha_{t}^{2}(\frac{\dot{\beta}_{t}}{\beta_{t}}-\frac{\dot{\alpha}_{t}}{\alpha_{t}})+\epsilon_{t}\right)^{2}\|\nabla\log\rho_{t}(x)-\hat{s}_{t}(x)\|^{2}_{2}\rho_{t}(x){\rm d}t{\rm d}x\,.$
@@ 2.4. Equivalence over scalar schedules
In [kingma2021variational], it was pointed out that in diffusion models, different noise schedules lead to the same variational lower bound. In the continuous setting, this corresponds to the KL divergence in path space. Our results generalize their discussion to stochastic interpolants and incorporate the step of a posteriori tuning of diffusion coefficients. $\lozenge$
Proposition 2.5 shows that the optimal KL accuracy in path space depends solely on the estimation of $\nabla\log q_{r}(x_{1}+rz)$: that is, from the perspective of KL divergence in path space, all linear scalar interpolants with independently coupled endpoints and one endpoint being Gaussian are statistically indistinguishable. This indicates that other metrics need to be explored if we want to select models for improved statistical efficiency. On the other hand, using nonlinear or matrix-valued instead of scalar schedules may potentially lead to different statistical efficiency, a direction of interest in future work.
@@ 3. Numerical Design by Optimizing Averaged Squared Lipschitzness
The discussion in the previous section does not consider numerical efficiency: while different scalar schedules are statistically equivalent, they lead to ODEs or SDEs with dramatically different regularity properties of the drift term. In this section, we explore how to choose interpolation schedules that enhance numerical efficiency. We focus on ODEs rather than SDEs for simplicity, noting that ODEs typically achieve better empirical performance due to their greater ease of integration [karras2022elucidating, dockhorn2022genie].
@@ 3.1. From one schedule to another
The proposition implies that we can easily change the interpolation schedule from one to another if we know the true drift functions. This also applies to the estimators of the drift functions, so we can tune the schedule at inference time rather than during training. Similar statements have appeared in the literature [kingma2021variational, karras2022elucidating]. We will use this fact in numerical experiments in Section 4. The natural question now is which schedule to choose in practice.
@@ 3.2. Optimizing averaged squared Lipschitzness
The time-dilated schedule studied in [aranguri2025optimizing] also resolve the mode missing issue:
@@ 3.5. High dimensional examples
This can be proved using the Cramér–Rao and Brascamp–Lieb inequalities; see [gao2024gaussian]. Therefore, similar to the Gaussian case, we can choose $\lambda^{\star}=\lambda_{M}^{-1}$. Then, with the schedule
@@ 4. Numerical Demonstrations
In this section, we conduct numerical experiments to demonstrate the improved numerical efficiency of the interpolation schedules designed in the previous section. We consider four classes of target distributions of increasing complexity: high-dimensional Gaussian random fields, high-dimensional Gaussian mixtures, invariant distributions of the stochastic Allen–Cahn equation, and invariant distributions of the stochastic Navier–Stokes equations. These examples range from exactly Gaussian to strongly non-Gaussian and serve to illustrate the robustness of the proposed schedule design principles.
In all experiments, the generative process is formulated as an ODE derived from stochastic interpolants. Except for the Gaussian and Gaussian mixture cases, where the drift fields are available in closed form, we parameterize the drift using the UNet architecture introduced in [ho2020denoising]. Training procedures, network architectures, and optimization hyperparameters follow [chen2024probabilistic] and are kept identical across different interpolation schedules. For numerical stability, the ODE is integrated from $t_{\min}=10^{-3}$ to $t_{\max}=1-10^{-3}$ in all experiments. Code to reproduce all results is available at https://github.com/yifanc96/GenerativeDynamics-NumericalDesign.git.
Sampling accuracy is evaluated using spectral diagnostics. For a generated sample $u$ (a one- or two-dimensional function in this paper), we compute the energy spectrum
$E(k)=\sum_{k\leq|m|_{2}<k+1}|\hat{u}(m)|^{2},$
where $\hat{u}(m)$ denotes the Fourier coefficients. In the Navier–Stokes case with vorticity formulation, this corresponds to the enstrophy spectrum. Reported spectra are obtained by averaging over a sufficiently large ensemble of independently generated samples.
@@ 4.1. Gaussians
We first consider a Gaussian random field with distribution $\mathsf{N}(0,\sigma^{2}(-\Delta+\tau^{2}\mathrm{I})^{-s})$ on $D=[0,1]^{2}$, where $\Delta$ denotes the Laplacian with homogeneous Dirichlet boundary conditions. We fix $s=3$, $\tau=1$, and $\sigma^{2}=(4\pi^{2}+\tau^{2})^{s}$, so that the field exhibits significant scale separation across Fourier modes. Target samples $x_{1}$ are drawn exactly from this distribution.
The noise variable $z$ in the stochastic interpolant is sampled from spatial white noise, corresponding to the Gaussian random field with $s=0$ and unit variance. The field is discretized on a uniform $N\times N$ grid, and the resulting generative ODE is integrated using a fixed-step fourth-order Runge–Kutta (RK4) method.
We compare the standard linear schedule $\beta_{t}=t$ with the designed schedule (3.11), which minimizes avg-Lip2 for Gaussian targets. Figure 2 shows representative samples at resolution $N=128$ generated using 20 RK4 steps. The designed schedule produces visibly smoother and more coherent samples. The right panel illustrates the schedules themselves, highlighting the rapid initial growth of the designed schedule, which reflects the fast speed required at early times.
Figure 3 compares the energy spectra of the true distribution and generated samples across resolutions. The designed schedule yields significantly more accurate spectra and maintains accuracy as $N$ increases, whereas the linear schedule deteriorates under refinement, reflecting its poorer numerical conditioning.
@@ 4.2. Gaussian mixtures
We next consider the high-dimensional Gaussian mixture distribution defined in (3.13) with dimension $d=1000$, mixture weight $p=0.3$, and mean vector $r=(1,1,\ldots,1)\in\mathbb{R}^{d}$. This distribution is strongly non-Gaussian and exhibits well-separated modes. The noise variable $z$ is sampled from $\mathsf{N}(0,\mathrm{I})$.
We compare the linear schedule $\beta_{t}=t$ with the approximate min-avg-Lip2 schedule (3.14), taking $\alpha_{t}=\sqrt{1-\beta_{t}^{2}}$ in both cases. The drift field is given analytically by Example 3.12, allowing us to isolate the numerical effects of the schedule without learning error. The ODE is integrated using only 2, 3, or 4 RK4 steps, with $10^{4}$ independent samples generated for each setting.
To assess mode recovery, we project the generated samples onto one dimension using PCA and fit a one-dimensional bimodal Gaussian mixture model. Table 1 reports the estimated smaller mixture weight. The linear schedule frequently collapses to a single mode when using few steps, while the approximate min-avg-Lip2 schedule accurately recovers both modes even under extremely coarse discretization.